# Theory of Semi-Deterministic Quantum Dot Placement in Heteroepitaxy

### v2

@@ Abstract
$V_{\mathrm{nc}}(\mathbf{r}-\mathbf{r}_{0},t)=\begin{cases}V_{\mathrm{seed}}(\mathbf{r}-\mathbf{r}_{0},t),&\text{if }|t-t_{0}|\leq\tau,\\[10.00002pt] V_{\mathrm{QD}}(\mathbf{r}-\mathbf{r}_{0}),&\text{if }\hat{n}(\mathbf{r}_{0},t)\geq n_{\mathrm{crit}},\\[10.00002pt] 0,&\text{otherwise (dissolution)},\end{cases}$ (13)
where the seed lifetime $\tau$ is drawn from an exponential distribution, $\tau\sim\exp(1/\tau_{c})$, and $t_{0}$ denotes the first-passage time for seeding, defined as,
$t_{0}=\min\left\{t\;\middle|\;\hat{n}(\mathbf{r}_{0},t)\geq n_{\mathrm{seed}}\right\}.$ (14)
@@
If no QD nucleation occurs within the interval $[t_{0},t_{0}+\tau]$, the potential is reset and the seeding process restarts at position $\mathbf{r}$. The system energy is updated by incorporating the time-dependent nucleation potential,
[…]
and the Langevin dynamics is modified accordingly.
@@
Without loss of generality, we adopt the following interaction forms in the simulations presented in this work. The boundary field is modeled as a delta-function sink potential, modulated by the (dimensionless) local strain profile $\sigma[\mathbf{r}_{\mathrm{bd}}(u)]$, which prevents adatoms from escaping once they reach the boundary,
@@
The pairwise repulsion between adatoms, $w$, and the nucleation potentials, $V_{\mathrm{nc}}$, are modeled using a screened power-law (Yukawa-like) potential $V$, expressed as,
$V\big{(}\mathbf{r}-\mathbf{r}^{\prime}\big{)}=\frac{\exp\left[-\left|\sqrt{2\,\mathbf{D}/\lambda}\,(\mathbf{r}-\mathbf{r}^{\prime})\right|\right]}{\left|\sqrt{2\,\mathbf{D}/\lambda}\,(\mathbf{r}-\mathbf{r}^{\prime})\right|^{\alpha}},$ (17)
where $\lambda$ is the finite adatom lifetime, associated with a constant evaporation rate. The power-law exponent $\alpha$ is microscopic parameter that characterize the strength and range of the many-body interactions.
IV \@slowromancapiv@. Linear Perturbation Theory in the diluted limit
In the early stage of the heteroepitaxy, thermally driven stochastic migration of adatoms leads to preferential deposition toward boundary edges that are orthogonal to the fast diffusion axis $\mathbf{\hat{x}}\equiv[1\bar{1}0]$, followed by lateral spreading along the boundary, as demonstrated numerically in Fig. 2C. When the average empirical density $\hat{n}$ over the exposed region $\Omega$ remains much smaller than the homogeneous density limit $\hat{n}_{\mathrm{hm}}=\Omega/\lambda_{w}^{2}$, the adatom dynamics are primarily governed by anisotropic diffusion. We refer to this regime as the diluted limit, where the exposed region $\Omega$ remains effectively adatom-free.
@@
We coarse-grain the empirical field $\hat{n}$ into a continuous density field $n$, and introduce a decay term to account for the finite lifetime of adatoms, characterized by a constant evaporation rate $\lambda$. Expanding locally around a boundary point $\mathbf{r}_{0,\mathrm{bd}}=(x_{0},y_{0})$, we assume that the convection–reaction–diffusion dynamics are separable to linear order along the principal axes. This yields a local density field $n_{0}$, modulated by separable spatial factors $f(x-x_{0})$ and $g(y-y_{0})$, such that,
@@
Expanded in Cartesian components around the boundary point $\mathbf{r}_{0,\mathrm{bd}}$, the local convection–diffusion–reaction equations take the form
@@
$\frac{n(u)}{n_{0}}=\int\mathcal{D}[\mathbf{c},\mathbf{c^{\prime}}]\exp\Big{[}$ $\int_{0}^{u}d\tilde{u}(\kappa(\tilde{u})+\mathbf{c}(\tilde{u})\,\mathcal{K}_{x}(\tilde{u})$ (27)
$+\mathbf{c^{\prime}}(\tilde{u})\,\mathcal{K}_{y}(\tilde{u}))\Big{]},$
@@
Figure 4: Geometric tuning of empirical adatom density fields under isotropic diffusion ($D_{x}=D_{y}$). A: By incorporating a nonzero repulsive pairwise interaction potential $w$ between adatoms, the resulting empirical adatom density field exhibits a smeared-out asymptotic density profile along the boundary contour. B: Starting from a lozenge boundary, elongation of the right-side flat edge with an aspect ratio $a=\alpha R$ induces local anisotropic confinement, leading to enhanced adatom accumulation near the elongated region. This demonstrates that local curvature can modulate the spatial profile of the adatom density field along the boundary. C: Simulations with various boundary geometries under initially uniform atomic deposition illustrate the flexibility and adaptability of the modeling framework in capturing geometry-tunable adatom density profiles and QD nucleation behavior along the boundary contour.
@@
In Figs. 2, 3, and 4, we systematically investigate QD nucleation profiles in the diluted limit across various boundary geometries, where stochastic seeding events occur exclusively at the boundary. The seeding probability is modulated by the adatom density field, which reflects the combined influences of biaxial diffusion, boundary field effects, and curvature. In the simulations, the ensemble-averaged empirical adatom density field $\langle\hat{n}\rangle$ is obtained from Langevin-based Monte Carlo sampling.
As shown in Fig. 2, this comparison is performed across multiple concentric circular boundaries of varying radii, under both isotropic and anisotropic diffusion conditions (Fig. 2B and 2C). When the QD nucleation potential $V_{\mathrm{nc}}$ is included, QDs are distributed along the boundary with spatial profiles that closely resemble the empirical adatom density fields. The spacing between boundary QDs is determined by the interplay of diffusion, geometry, and $V_{\mathrm{nc}}$. For instance, once a QD nucleates at a given location along the boundary, it increases the local chemical potential, generating effective repulsive interactions that redirect subsequent adatoms to other regions.
@@
V \@slowromancapv@. secondary QD formation on pristine surface
@@
A stable InAs/GaAs QD on a pristine surface has a typical size range of 20 to 25 nm, and QD nucleation sites are highly sensitive to material properties and growth conditions. The interaction length scale $\ell_{\mathrm{QD}}$ of the nucleation potential $V_{\mathrm{nc}}$ can extend to several tens of nanometers, or even up to hundreds of nanometers, as inferred from the observed relationship between QD density and growth rate [47, 48]. In contrast, stable QDs formed at the boundary often exhibit irregular shapes and stress profiles, which can result in comparable or even greater interaction length scales, ranging from a few nanometers to several hundred nanometers. The repulsive field generated by those stable QDs at the boundary not only affects the migration of nearby adatoms but also interacts with adatom densities in the bulk, influencing the global spatial accumulation and redistribution of newly deposited adatoms, which can be used to statistically control the seeding location on the pristine bulk surface. This concept is illustrated schematically in Fig. 5.
@@
As demonstrated in Fig. 6A, to model boundary-mediated secondary QD nucleation, we begin with a elongated lozenge boundary field design and vary the right-side diagonal coordinate $(a,0)$ from $-0.6L$ to $2L$, while conserving the total bulk area. The epitaxial deposition of new atoms is introduced sequentially at a rate much slower than the time required for adatoms to reach a steady-state distribution through Langevin dynamics. We then record the spatial distribution corresponding to the first QD nucleation site, as illustrated in the middle panel in Fig. 5. This process is repeated 400 times using different initial random adatom configurations (uniformly distributed). Depending on the value of $a$, as shown in Fig. 6B and Fig. 7, the resulting spatial distribution of stable QDs exhibits a boundary-dependent profile. The observed inhomogeneity in the QD distribution reveals statistically preferential nucleation sites. As detailed in Appendix D Fig. 8, we further perform a full parameter sweep over $a\in[-0.6L,2L]$ for a hexagonal boundary field with a $60^{\mathrm{o}}$ orientation. While the microscopic parameters of the Langevin dynamics remain unchanged, the hexagonal boundary field produces a more spatially extended adatom density field compared to the elongated lozenge boundary field shown in Fig. 7.
VI \@slowromancapvi@. Discussion and Conclusion
@@
In this work, we have theoretically and numerically investigated the interplay between geometry, diffusion, and stochastic dynamics in the context of quantum dot (QD) nucleation from empirical adatom density fields. Our results illustrate how geometric properties of boundary contours, such as curvature and orientation relative to anisotropic diffusion axes, along with intrinsic microscopic properties, such as adatom interactions and QD nucleation dynamics, shape the spatial patterns of QD formation.
@@
This work enables the deterministic positioning of high optical quality self-assembled quantum dots on pristine surfaces, opening a new direction for next-generation on-chip photonic devices.
@@ Appendix B Appendix B: Path Integral Formulation and Mean-field Approximation
$=\prod_{i=0}^{N}\sum_{c,c^{\prime}=\pm 1}A^{i}_{c,c^{\prime}}\exp\left(du_{i}\,k_{x}^{c}|T_{x}(u_{i})|+k_{y}^{c^{\prime}}|T_{y}(u_{i})|\right).$

### v4

@@ Abstract
$V_{\mathrm{nc}}(\mathbf{r}-\mathbf{r}_{0},t)=\begin{cases}V_{\mathrm{seed}}(\mathbf{r}-\mathbf{r}_{0},t),&\text{if }|t-t_{0}|\leq\tau,\\[10.00002pt] V_{\mathrm{QD}}(\mathbf{r}-\mathbf{r}_{0}),&\text{if }\hat{n}(\mathbf{r}_{0},t)\geq\hat{n}_{\mathrm{self-limited}},\\[10.00002pt] 0,&\text{otherwise (dissolution)},\end{cases}$ (16)
To model regimes A and B, we introduce a localized seed potential $V_{\mathrm{seed}}$ at $\mathbf{r}_{0}$ whenever the empirical density exceeds a subcritical threshold $\hat{n}_{\mathrm{seed}}$. The seeding potential serves as a binding potential $V_{\mathrm{seed}}$ that anchors nearby adatoms, promoting local adatom accumulation and cluster formation.
We model $V_{\mathrm{seed}}$ as a radially symmetric trapping potential with sufficient depth (relative to thermal fluctuations), and the effective area of this potential is taken to be the underlying Monte Carlo grid size $d\Omega$. This seed persists for a finite duration $\tau_{c}$, during which the local seed is allowed to either evolve toward the QD self-limited size with density $\hat{n}_{\mathrm{self-limited}}$ at $\mathbf{r}_{0}$, corresponding to the second minimum in $\Delta G$, or dissolve back to the uniform state (no clusters). The seed lifetime $\tau$ is drawn from an exponential distribution, $\tau\sim\exp(1/\tau_{c})$, and $t_{0}$ denotes the first-passage time for seeding, defined as,
$t_{0}=\min\left\{t\;\middle|\;\hat{n}(\mathbf{r}_{0},t)\geq\hat{n}_{\mathrm{seed}}\right\}.$ (17)
If no QD nucleation occurs within the interval $t_{0},t_{0}+\tau]$, the seed potential is removed, corresponding to dissolution. Both the subcritical threshold $\hat{n}_{\mathrm{seed}}$ and supercritical density $\hat{n}_{\mathrm{supercritical}}$ determined by detailed growth conditions. Since reaching the supercritical density $\hat{n}_{\mathrm{supercritical}}$ triggers coherent QD growth toward the self-limited density $\hat{n}_{\mathrm{self-limited}}$, for computational simplicity we merge these two points into a single reference and denote it as the critical density $\hat{n}_{\mathrm{crit}}$ (see Fig. 2). The supercritical and critical densities are taken to be of the same order of magnitude as the homogeneous background state, $\hat{n}_{\mathrm{crit}}\approx 10^{3}\hat{n}_{\mathrm{seed}}$ and $\hat{n}_{\mathrm{seed}}\approx 10^{2}\hat{n}(\mathbf{r}_{0},0)$.
As the empirical density reaches $\hat{n}_{\mathrm{crit}}$ within the time window $\tau_{c}$, the seed potential is irreversibly converted into a stable QD potential $V_{\mathrm{QD}}$, indicating successful traversal of the nucleation barrier into regime C, creating a permanent QD at $\mathbf{r}_{0}$ where it subsequently generate repulsion ($V_{\mathrm{QD}}$) to newly deposited adatoms, reflecting the stabilization of the dot and the inhibition of further growth beyond its self-limited size.
The detailed interatomic interactions among adatoms, between adatoms and QDs, and among QDs depend on the growth conditions and specific atomic properties, such as bond ionicity and van der Waals interactions, and are typically obtained from empirical studies [44]. For the purpose of demonstrating semi-deterministic QD placement, we model the pairwise repulsion between adatoms (dipole–dipole) and between adatoms and QDs (dipole–monopole) [60], denoted $V_{\mathrm{nc}}$, using a generic screened Coulomb-like potential $V$, expressed as,
$V\big(\mathbf{r}-\mathbf{r}^{\prime}\big)\sim\frac{\exp\left[-\xi\left|\boldsymbol{\ell}^{-1}(\mathbf{r}-\mathbf{r}^{\prime})\right|\right]}{\left|\boldsymbol{\ell}^{-1}(\mathbf{r}-\mathbf{r}^{\prime})\right|^{\alpha}}.$ (18)
Here, $\xi$ is a screening parameter that sets the scale, and the characteristic decay length-scale tensor is defined as $\boldsymbol{\ell}=\sqrt{\mathbf{D}/\lambda}$, where $\lambda$ is the constant desorption rate that sets the interaction scale. The power-law exponent $\alpha$ is a microscopic parameter that characterizes the type and range of the many-body interactions. For example, a dipole–monopole interaction corresponds to $\alpha=2$, and a dipole–dipole interaction corresponds to $\alpha=3$.
@@
The system energy is updated by incorporating the time-dependent nucleation potential,
$U_{\mathrm{sys}}\to U_{\mathrm{sys}}+\frac{1}{2}\iint_{\Omega}\hat{n}(\mathbf{r},t)\,V_{\mathrm{nc}}(\mathbf{r}-\mathbf{r}_{0},t)\,\hat{n}(\mathbf{r}_{0},t)\,d\mathbf{r}\,d\mathbf{r}_{0},$ (19)
and the Langevin dynamics is modified accordingly. The pairwise interaction $w$ in $U_{\mathrm{sys}}$ (Eq. 8) is approximated as a point-like elastic dipole interaction at long range and is assumed to be repulsive. This discourages adatoms from clustering too early, favors nucleation at distinct sites, and contributes to the size uniformity of QDs [44, 60].
@@
The boundary field is modeled as a delta-function sink potential, modulated by the (dimensionless) local strain profile $\sigma[\mathbf{r}_{\mathrm{bd}}(u)]$, which prevents adatoms from escaping once they reach the boundary,
@@
IV IV. Linear Perturbation Theory in the diluted limit
In the early stage of heteroepitaxy, thermally driven stochastic migration of adatoms causes preferential deposition at boundary edges orthogonal to the fast diffusion axis $\mathbf{\hat{x}}\equiv[1\bar{1}0]$ [27], followed by lateral spreading along the boundary, as shown numerically in Fig. 3C. When the average empirical density $\hat{n}$ over the exposed region $\Omega$ remains much smaller than the homogeneous density limit, the adatom dynamics are primarily governed by anisotropic diffusion. We refer to this regime as the diluted limit, where the exposed region $\Omega$ remains effectively adatom-free.
@@
We coarse-grain the empirical field $\hat{n}$ into a continuous density field $n$, and introduce a decay term to account for the finite lifetime of adatoms, characterized by a constant desorption rate $\lambda$. Expanding locally around a boundary point $\mathbf{r}_{0,\mathrm{bd}}=(x_{0},y_{0})$, we assume that the convection–reaction–diffusion dynamics are separable to linear order along the principal axes. This yields a local density field $n_{0}$, modulated by separable spatial factors $f(x-x_{0})$ and $g(y-y_{0})$, such that,
@@
Expanded in Cartesian components around the boundary point $\mathbf{r}_{0,\mathrm{bd}}$, the local convection–diffusion–reaction equations take the form,
@@
[…]
$)],\displaystyle+\mathbf{c^{\prime}}(\tilde{u})\,\mathcal{K}_{y}(\tilde{u}$$)\Big$],
@@
Figure 5: Geometric tuning of empirical adatom density fields under isotropic diffusion ($D_{x}=D_{y}$). A: By incorporating a nonzero repulsive pairwise interaction potential $w$ between adatoms, the resulting empirical adatom density field exhibits a smeared-out asymptotic density profile along the boundary contour. B: Starting from a lozenge boundary, elongation of the right-side flat edge with an aspect ratio $a/R$ induces local anisotropic confinement, leading to enhanced adatom accumulation near the elongated region. This demonstrates that local curvature can modulate the spatial profile of the adatom density field along the boundary. C: Simulations with various boundary geometries under initially uniform atomic deposition illustrate the flexibility and adaptability of the modeling framework in capturing geometry-tunable adatom density profiles and QD nucleation behavior along the boundary contour.
@@
In Figs. 3, 4, and 5, using Eq. 16, we systematically investigate QD nucleation profiles in the diluted limit across various boundary geometries, where stochastic seeding events occur almost exclusively at the boundary. The seeding probability is modulated by the adatom density field, which reflects the combined influences of biaxial diffusion, boundary field effects, and curvature. In the simulations, the ensemble-averaged empirical adatom density field $\langle\hat{n}\rangle$ is obtained from Langevin-based Monte Carlo sampling.
As shown in Fig. 3, this comparison is performed across multiple concentric circular boundaries of varying radii, under both isotropic and anisotropic diffusion conditions (Fig. 3B and 3C). When the QD nucleation potential $V_{\mathrm{nc}}$ is included, QDs are distributed along the boundary with spatial profiles that closely resemble the empirical adatom density fields. The spacing between boundary QDs is determined by the interplay of diffusion, geometry, and $V_{\mathrm{nc}}$. For instance, once a QD nucleates at a given location along the boundary, it increases the local chemical potential, generating effective repulsive interactions that redirect subsequent adatoms to other regions.
@@
As additional adatoms are deposited, they become “aware” of the surrounding strain field, generated by both the engineered boundary and primary QDs nucleated at high-curvature boundary regions. This evolving strain landscape dynamically alters subsequent QD nucleation sites at the boundaries and within the bulk (pristine surface). Building on this, in the following section we utilize strain sensitivity, together with engineered boundary fields, to guide QD nucleation in a controlled manner, enabling semi-deterministic QD placement.
V V. secondary QD formation on pristine surface
@@
A stable InAs/GaAs QD on a pristine surface has a typical size range of 20 to 25 nm with nucleation sites that are both highly sensitive to material properties and growth conditions. The interaction length scale of the nucleation potential $V_{\mathrm{nc}}$ can extend to several tens of nanometers, or even up to hundreds of nanometers, as inferred from the observed relationship between QD density and growth rate, or QD array periodicity  [62, 63]. In contrast, stable QDs formed at the boundary often exhibit irregular shapes and accumulated stress profiles, which can result in comparable or even greater interaction length scales, ranging from a few nanometers to several hundred nanometers. The repulsive field generated by those stable QDs at the boundary not only affects the migration of nearby adatoms but also interacts with adatom densities in the bulk (pristine surface), influencing the global spatial accumulation and redistribution of newly deposited adatoms, which can be used to statistically control the seeding location on the pristine bulk surface. This concept is illustrated schematically in Fig. 6.
@@
As demonstrated in Fig. 7A, to model boundary-mediated secondary QD nucleation, we begin with a elongated lozenge boundary field design and vary the right-side diagonal coordinate $(a,0)$ from 0.6L-0.6L to $2L$, while conserving the total bulk area. The epitaxial deposition of new atoms is introduced sequentially at a rate much slower than the time required for adatoms to reach a steady-state distribution through Langevin dynamics. We then record the spatial distribution corresponding to the first QD nucleation site, as illustrated on the far right side of Fig. 6E. This process is repeated 400 times using different initial random adatom configurations (uniformly distributed). Depending on the value of $a$, as shown in Fig. 7B and Fig. 8, the resulting spatial distribution of stable QDs exhibits a boundary-dependent profile. The observed inhomogeneity in the QD distribution reveals statistically preferential nucleation sites. As detailed in Appendix D Fig. 9, we further perform a full parameter sweep over $a\in[-0.6L,2L]$ for a hexagonal boundary field with a $60^{\mathrm{o}}$ orientation. While the microscopic parameters of the Langevin dynamics remain unchanged, the hexagonal boundary field produces a more spatially extended adatom density field compared to the elongated lozenge boundary field shown in Fig. 8.
Intuitively, semi-deterministic placement is expected: upon achieving primary QD distribution along the engineered boundary field, subsequent adatoms mainly diffuse along the fast axis (and hardly along the slow axis) and experience secondary interactions that can be significantly slowed down when near primary QDs, leading to adatom accumulation over time at those locations.
While the above simulation is done at mild anisotropy, $D_{x}=3D_{y}$, for illustration purposes, in the case of the GaAs(001)-$\beta_{2}(2\times 4)$ surface with diffusion axes $\mathbf{\hat{x}}\equiv[1\bar{1}0]$ and $\mathbf{\hat{y}}\equiv[110]$, the diffusion constant and anisotropy, $D_{x}\sim 10^{3}D_{y}\sim 1\times 10^{-3}~\mu\mathrm{m}^{2}/\mathrm{s}$ along the two directions, follow at typical temperatures $T\sim 500~\mathrm{K}$ [46], which gives the relative diffusion length scale $\ell_{x}/\ell_{y}\sim 32$. This suggests that an efficient redistribution of the empirical density field occurs along the fast diffusion axis. Therefore, we predict that by engineering boundary fields with anisotropy along the fast diffusion axis (for example, shape edges across two opposite boundaries, see the comparison in Appendix D, Fig. 8 and Fig. 9), the modulation effects from both the boundary fields and the on-boundary primary nucleated QDs would exert the strongest influence on subsequently deposited adatoms, leading to more pronounced modulation that enables semi-deterministic placement of secondary QDs.
Since we assume a quasi-stationary limit, the deposition rate must be sufficiently low to allow adatoms to reach quasi-equilibrium. Let’s consider a unit deposition time, with adatom diffusion lengths of $\ell_{x}\sim 32\,\mathrm{nm}$ and $\ell_{y}\sim 1\,\mathrm{nm}$. The effective diffusion area is then $d\Omega\sim\ell_{x}\cdot\ell_{y}=32\,\mathrm{nm}^{2}$. The deposition flux $F$ (atoms per unit area per unit time) corresponding to $N$ atoms deposited in this area is given by $F\sim N/d\Omega$. In the quasi-stationary limit, within a unit time, the number of deposited adatoms over the effective area should be much smaller than one adatom per diffusion area. Taking the bound $N=1$, the corresponding flux is $F=1/(32\,\mathrm{nm}^{2}\cdot 1\,\mathrm{s})\approx 3.125\times 10^{4}\,\mathrm{atoms/(\mu m^{2}\cdot s)}$. To convert this to a deposition rate in monolayers per second (ML/s), we use the lattice constant of GaAs ($a=0.565\,\mathrm{nm}$) to calculate the surface atomic density as $n_{s}=2/a^{2}\sim 6.26\times 10^{6}\,\mathrm{atoms/\mu m^{2}}$. This gives the upper bound on the deposition rate $R_{\mathrm{bound}}=F/n_{s}\sim 5\times 10^{-3}\,\mathrm{ML/s}$, which is consistent with typical growth rates used for InAs QDs on GaAs(001) substrates [61]. To achieve the quasi-stationary limit, the empirical deposition rate must satisfy $R\ll R_{\mathrm{bound}}$ around temperature $T\sim 500$ K, which can be realized under experimental conditions.
Practically, designing such boundary fields is straightforward with standard fabrication capabilities. A typical hard mask can first be uniformly deposited across the substrate (e.g., a (001)-oriented GaAs surface), and then selectively etched to re-expose patterned regions for regrowth.
VI VI. Discussion and Conclusion
@@
Over the past few decades, strain-mediated interactions between QDs and adatoms have been extensively studied theoretically, focusing on QD morphology and phase transitions, yet few studies have explored their use in practical applications such as QD placements. In this work, we have theoretically and numerically investigated the interplay between geometry, diffusion, and stochastic dynamics in the context of quantum dot (QD) nucleation from empirical adatom density fields. Our results illustrate how geometric properties of boundary contours, such as curvature and orientation relative to anisotropic diffusion axes, along with intrinsic microscopic properties, such as adatom interactions and QD nucleation dynamics, shape the spatial patterns of QD formation.
@@
This work enables the semi-deterministic positioning of high optical quality self-assembled quantum dots on pristine surfaces, opening a new direction for next-generation on-chip photonic devices.
@@ Appendix B Appendix B: Path Integral Formulation and Mean-field Approximation
$=\prod_{i=0}^{N}\sum_{c,c^{\prime}=\pm 1}A^{i}_{c,c^{\prime}}\exp\left(du_{i}(\,k_{x}^{c}|T_{x}(u_{i})|+k_{y}^{c^{\prime}}|T_{y}(u_{i})|)\right).$